# Understanding the Performance Gap in Preference Learning: A Dichotomy of RLHF and DPO

### v1

@@ Appendix D Implementation Details
Hyper-parameters. The maximum length is set as $256$. The prompt template is “BEGINNING OF CONVERSATION: USER: [prompt] ASSISTANT: [response]”. SFT: The hyper-parameter setting is based on Dong et al. [2024]. We use a batch size $32$. Online training: The hyper-parameter setting is based on Dong et al. [2024]. We use a batch size $32$, a learning rate $5\text{e}-7$, and a gradient accumulation step $2$. We train for $3$ iterations, each for $2$ epochs. We set $r_{\textup{margin}}=0.4,1,4$ for verifications of Condition 1, and set $r_{\textup{margin}}=1$ for verifications of Conditions 2, 3 and 4. Offline training: The hyper-parameter setting is based on Zhou et al. [2024]. We use a batch size $4$, a learning rate $1\text{e}-4$, and a gradient accumulation step $2$. We train for $3$ epochs (when training reward model on $9$k data of PKU-SafeRLHF-safer, we train $6$ epochs for higher training accuracy). We haven’t extensively tuned these hyper-parameters.
Computation resources. Our experiments are conducted on NVIDIA RTX A6000. SFT and Online training: We adopt $4$ workers, each taking up $35,000$M of memory, running for $2$-$3$ hours. Offline training: We adopt $1$ worker, which takes up $25,000$M of memory and runs for up to $40$ minutes.

### v2

@@ Appendix C Implementation Details
Hyper-parameters. The maximum length is set as $256$. The prompt template is “BEGINNING OF CONVERSATION: USER: [prompt] ASSISTANT: [response]”. SFT: The hyper-parameter setting is based on Dong et al. [2024]. We use a batch size $32$. Online training: The hyper-parameter setting is based on Dong et al. [2024]. We use a batch size $32$, a learning rate $5\text{e}-7$, and a gradient accumulation step $2$. We train for $3$ iterations, each for $2$ epochs. We set $r_{\textup{margin}}=0.4,1,4$ for verifications of Condition 1, and set $r_{\textup{margin}}=1$ for verifications of Conditions 2, 3 and 4. Offline training: The hyper-parameter setting is based on Zhou et al. [2024]. We use a batch size $4$, a learning rate $1\text{e}-4$, and a gradient accumulation step $2$. We train for $3$ epochs (when training reward model on 99k data of PKU-SafeRLHF-safer, we train $6$ epochs for higher training accuracy). We haven’t extensively tuned these hyper-parameters.
Computation resources. Our experiments are conducted on NVIDIA RTX A6000. SFT and Online training: We adopt $4$ workers, each taking up 35,00035, [^0] M of memory, running for 22-33 hours. Offline training: We adopt $1$ worker, which takes up 25,00025, [^0] M of memory and runs for up to $40$ minutes.